OpenStudy (a1234):

θ = 16 degs y = 17 find the exact value of r.

OpenStudy (a1234):

 Is r = 17/cos16 degs

OpenStudy (johnweldon1993):

\[\large cos(\theta) = \frac{y}{r}\] \[\large r = \frac{y}{cos(\theta)}\] Good job